# An Adjoint Formulation of Energetic Particle Confinement

### v1

@@ Abstract
An adjoint formulation of energetic particle confinement in axisymmetric geometry is derived and evaluated using a Physics-Informed Neural Network (PINN). The PINN estimates the escape time of energetic ions by solving an inhomogeneous adjoint of the drift kinetic equation with a Lorentz collision operator, yielding predictions of the escape time of fast ions in tokamak geometry due to direct ion orbit loss and collisional transport. This is the first time a PINN has been used to solve the drift kinetic equation in tokamak geometry, a challenging problem due to the large time scale separation present between the rapid transit time of energetic ions, and their slow collision time scale. It is shown that a careful and intentional design of a PINN is able to learn the escape time for the majority of the geometry considered, suggesting a path toward constructing a rapid surrogate for use in a broader optimization framework.
@@ I Introduction
In the present work, an initial step toward a framework through which the confinement time of an energetic ion or electron can be efficiently estimated across a broad range of parameter regimes is developed. While evaluating the confinement time of an energetic particle is often of greatest importance for non-axisymmetric magnetic fields arising from error fields Goldston et al. (1981), magnetic islands Carolipio et al. (2002) or energetic particle modes Heidbrink and White (2020) in a tokamak, or for a 3D stellarator equilibrium Galeev et al. (1969), this work will consider the simpler problem of an axisymmetric tokamak equilibrium. Our motivation for considering this simpler problem is largely to investigate the potential for evaluating metrics of energetic particle transport via the solution to an inhomogeneous adjoint problem. As described below, the inhomogeneous adjoint of the drift kinetic equation provides a direct means of computing the average confinement time of an ion or electron as a function of the particle’s initial phase space position Särkimäki (2020), a quantity we will refer to as the escape time. While this adjoint problem can be solved by a variety of numerical methods, we will consider a Physics-Informed Neural Network (PINN) Raissi et al. (2019) in the present work. Physics-Informed Neural Networks correspond to a machine learning framework capable of integrating both physics and data into the training of a neural network. A distinguishing feature of a PINN is that it may be extended to learn the parametric dependence of the solution of a PDE Sun et al. (2020). Noting that optimizing the transport properties of energetic particle orbits often requires exploring a broad range of magnetic field configurations, obtaining a rapid surrogate capable of computing quantities of interest such as the escape time of energetic particles provides a powerful tool for assessing energetic particle transport.
@@
Our aim in the present paper is not to provide such a parametric solution, but to instead investigate the ability of a PINN to solve the drift kinetic equation for an energetic ion population. Specifically, while PINNs have previously been used to solve the relativistic Arnaud et al. (2024); McDevitt et al. (2025); Arnaud et al. (2025) and non-relativistic McDevitt and Tang (2024) Fokker-Planck equations, the large time scale separation between the rapid bounce or transit times and slow collision time of energetic particles in tokamak geometry, represents a far more challenging problem. Utilizing advanced optimization algorithms Urbán et al. (2025); Kiyani et al. (2025); Wang et al. (2025), we aim to show that PINNs are able to robustly resolve several features of this challenging system, even in the absence of data. As described further below, as the energy of the ions is increased, leading to a greater time scale separation between the fast bounce/transit motion and slow collision time of energetic ions, our implementation of a PINN struggles to quantitatively describe the escape time of the best confined ion orbits, though it does resolve the solution in the outer region of the tokamak. In future work, we will seek to overcome this limitation by directly utilizing small quantities of Monte Carlo data generated by a traditional particle based drift kinetic solver.
@@
As an alternative means of evaluating the fast ion escape time, we introduce a new drift-kinetic solver based on the JAX framework. While in the present work this solver is only used to verify solutions from the PINN, future work will leverage particle data generated by this drift-kinetic solver to help accurately resolve the escape time of well confined ion orbits, along with facilitating the training of the PINN over a broad range of plasma parameters. This extended PINN framework, trained using both physics and particle data, would thus provide an efficient surrogate for fast ion confinement appropriate for use inside optimization loops.
@@
The remainder of this paper is organized as follows. Section II derives an adjoint problem that will be solved in this paper. A brief overview of the physics-informed deep learning framework and JAX particle-based solver are described in Sec. III. The escape time of an energetic ion species is computed in Sec. IV. Conclusions and a discussion of future directions are given in Sec. V
@@ II Adjoint of Steady State Drift Kinetic Equation
In this section we will define an adjoint problem of the inhomogeneous steady state drift kinetic equation. Our starting point will be to define the Green’s function $F_{s}\left(\mathbf{z};\mathbf{z}_{0}\right)$ for the steady state adjoint equation Helander and Sigmar (2002):
@@
$C_{s}\left(F_{s}\right)=\frac{\nu_{D}}{2}\frac{\partial}{\partial\xi}\left[\left(1-\xi^{2}\right)\frac{\partial F}{\partial\xi}\right]{\color[rgb]{1,0,0},}$ (2)
@@
where $r_{wall}$ is the location of the surrounding wall, taken as the plasma minor radius $a$. With these boundary conditions no flow of particles is allowed to enter the spatial domain, but particles are free to exit. The resulting Green’s function can thus be physically interpreted as the steady state particle distribution resulting from a unit source located at $\mathbf{z}=\mathbf{z}_{0}$.
To derive an adjoint problem for the escape time of an energetic particle, we begin by multiplying the left hand side of Eq. (1a) by the solution to the adjoint equation $T_{s}$, which after successive integrations by parts yields:
@@
$=\int d^{3}vd^{3}xF_{s}\left[-\mathbf{\dot{X}}\cdot\nabla T_{s}-\mathbf{\dot{X}}\frac{\partial T_{s}}{\partial v_{\|}}-C^{*}_{s}\left(T_{s}\right)\right]$
$+\int d^{3}v\int d\Gamma\left.\left[U_{r}T_{s}F_{s}\right]\right|_{r=a}.$ (5)
@@
An adjoint problem for the escape time of an energetic particle can be derived by considering the inhomogeneous adjoint equation Liu et al. (2017)
@@
Substituting Eq. (6) along with the steady state Green’s function defined by Eq. (1a) into the adjoint relation defined by Eq. (5), yields
@@
Further noting that $N_{s}$ will scale directly with the confinement time of the particles, $T_{s}\left(\mathbf{x}_{0},\mathbf{v}_{0}\right)$ will act as a measure of how well particles injected at a given phase space location are confined on average, a quantity we will refer to as the escape time.
@@ III Physics-Constrained Deep Learning Framework
The primary focus of the remainder of this work will be the solution of the adjoint problem derived in Sec. II. A brief overview of our implementation of a PINN is given in Sec. III.1, where the reader is referred to Refs. Cuomo et al. (2022); Wang et al. (2023); Toscano et al. (2025) for additional information on this rapidly expanding area. In addition, a newly developed GPU accelerated particle-based solver to the drift kinetic equation is described in Sec. III.2. For the present work, this particle-based solver will be used to validate the PINN’s solutions to the drift kinetic equation.
@@ III.1 Physics-Informed Neural Networks
Physics-constrained machine learning methods Karniadakis et al. (2021) seek to incorporate physical constraints into the training of a neural network. Physics-informed neural networks correspond to a prominent example of this area. An attractive feature of PINNs is that they can be used to solve PDEs and ODEs directly in the absence of data van Milligen et al. (1995); Raissi et al. (2019), or they provide a mechanism through which physical constraints can be used to enrich sparse data sets Cai et al. (2021); Mathews et al. (2022); McDevitt et al. (2024). In its simplest form, the loss function of a PINN can be written as:
@@
where $T$ is the quantity we seek to predict, the escape time for the current work, $\mathcal{R}$ is the residual of the governing PDE, $w_{PDE}$ is a scalar weight applied to the PDE bias of the loss, $\mathbf{z}=\left(\mathbf{x},\mathbf{v}\right)$ are the independent variables, and $\bm{\lambda}$ describes physical parameters, such as collisionality or inverse aspect ratio. The first term represents a bias due to data, where the data may be boundary conditions, or synthetic or experimental data. The second term is a bias due to the governing equations, which for the present work will be the inhomogeneous adjoint of the steady state drift kinetic equation. In the absence of synthetic or experimental data, the first term will only contain the boundary conditions. Successfully minimizing the loss will thus correspond to identifying a solution to the PDE while satisfying the boundary conditions. Unlike traditional PDE solvers, a parametric solution to a PDE can be straightforwardly obtained by training across a broad range of physics parameters $\bm{\lambda}$. Thus, while the training time of such a model may be substantially longer than a traditional PDE solver, once trained, the fast inference time of a PINN allows for fast online prediction time across a broad range of plasma conditions. This property can be used for the development of fast surrogate models, where some recent examples include runaway electron generation McDevitt (2023); Arnaud et al. (2025), MHD equilibrium Jang et al. (2024), plasma thrusters Luo et al. (2025), or non-thermal ion distributions McDevitt and Tang (2024).
A key element in obtaining accurate results from PINNs corresponds to (i) the identification of an appropriate neural network architecture, (ii) the phase space weighting of the residual, and (iii) the selection of an efficient optimization algorithm for minimizing the loss. Regarding item (i) we will employ a fully connected feedforward neural network, but with an ‘output layer’ that enforces positivity as a hard constraint. In so doing, this both prevents certain unphysical solutions from being predicted by the PINN, and perhaps more importantly, narrows the space of solutions that the optimizer searches for. Defining $T_{NN}$ to be the output of the hidden layers of the neural network, a simple transformation that enforces positivity is given by
@@
In addition to enforcing positivity, the exponential dependence of $T_{s}$ on the output of the hidden layers $T_{NN}$ provides a natural means of capturing the large disparity in escape times for energetic ions. Specifically, ions born near the plasma core are expected to be well confined, where they will only be lost by the relatively slow process of collisional transport, in contrast to those born near the edge, which may be lost on a very short timescale.
@@
$\text{Loss}=\frac{1}{N_{bdy}}\sum^{N_{bdy}}_{i=1}\left[T_{s,i}-T_{s}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\right]^{2}+\frac{w_{pde}}{N_{PDE}}\sum^{N_{PDE}}_{i=1}\left[\mathcal{F}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\mathcal{R}_{IH}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\right]^{2}.$ (13)
@@
Here, the boundary conditions defined by Eq. (8) are enforced by the first term, $\mathcal{R}_{IH}$ is the residual of the inhomogeneous adjoint defined by Eq. (6), and the choice of $A$ sets how much emphasis is given to the edge region. In particular, noting that near the edge $T_{s}$ will be of the order of a transit time, since ions on loss orbits will directly escape from the plasma, if $A$ is chosen to satisfy $A\ll 1$, Eq. (12) will scale as $\propto 1/T_{s}$, such that the edge region will be heavily weighted, but the inner region of the plasma with $T_{s}\gg 1$, will receive minimal weighting. We thus expect this choice to allow for the outer region of the plasma to be accurately described, at the expense of poor accuracy in the inner region. In contrast, by choosing a larger value of $A$, improved accuracy for the inner region can be achieved, at the expense of having a looser connection to the edge boundary condition, which is vital for achieving a physically meaningful solution. We have found that values of $A$ between 10 and 100 allow for robust convergence of the PINN.
In addition, the relative weight of the PDE residual compared to the boundary condition is set by the scalar $w_{PDE}$ weighting the first term in Eq. (13). We have found $w_{PDE}=100$ provides relatively rapid convergence of the loss. Finally, since we anticipate the edge region to contain the most rapid variation of the escape time, we will pack a disproportionate number of training points into this region.
@@
The last key component of our PINN implementation will be the choice of an effective optimization algorithm. While the ADAM optimizer Kingma and Ba (2014) is used ubiquitously across a range of machine learning applications, this first order optimization algorithm often is not able to achieve low losses for challenging PDEs. A common strategy is thus to use ADAM for the first phase of training, after which limited memory BFGS (L-BFGS) Liu and Nocedal (1989), a second order optimization strategy, is used to more tightly converge the loss. Motivated by Refs. Wang et al. (2025) and Urbán et al. (2025); Kiyani et al. (2025), we have adopted a different optimization strategy that we have found to yield substantially improved accuracy. The first optimizer that we apply corresponds to the quasi second order optimizer SOAP Vyas et al. (2024). This scalable optimizer exhibits similar computational performance as ADAM, but can yield losses of up to an order of magnitude smaller than those achievable by ADAM across a broad range of challenging applications of PINNs Wang et al. (2025). For the second stage of optimization, we utilize the quasi-Newton Self-Scaled Broyden (SSBroyden) method Al-Baali et al. (2014). Like L-BFGS, SSBroyden approximates the Hessian, but introduces additional degrees of freedom into the optimization algorithm that have yielded substantial improvements in training across a range of complex PDEs Urbán et al. (2025); Kiyani et al. (2025). While each training epoch is substantially more computationally demanding compared to ADAM or SOAP, we have found it results in substantially faster convergence of the loss for the adjoint problem treated in this work. The primary limitations of SSBroyden as an optimization routine is that the SciPy implementation used in the present work does not efficiently utilize GPUs or support mini-batching. These limitations substantially increase the training time of the SSBroyden phase of training, and limit the number of training points employed.

### v2

@@ Abstract
An adjoint formulation of energetic particle confinement in axisymmetric tokamak geometry is derived and evaluated using a physics-informed neural network (PINN). The PINN estimates the mean escape time of energetic ions by solving an inhomogeneous adjoint of the drift kinetic equation with a Lorentz collision operator, yielding predictions of fast ion loss in tokamak geometry due to direct ion orbit loss and collisional transport. To our knowledge, this is the first time a PINN has been used to solve the drift kinetic equation in tokamak geometry, a challenging problem due to the large time scale separation between the rapid transit time of energetic ions and their slow collisional time scale. It is shown that a careful and intentional design of a PINN is able to learn the mean escape time across the majority of the plasma volume, suggesting a path toward constructing a rapid surrogate for use within a broader optimization framework.
@@ I Introduction
In the present work, an initial step toward a framework through which the confinement time of an energetic ion or electron can be efficiently estimated is developed. While evaluating the confinement time of an energetic particle is often of greatest importance for non-axisymmetric magnetic fields arising from error fields Goldston et al. (1981), magnetic islands Carolipio et al. (2002) energetic particle modes Heidbrink and White (2020), or for a 3D stellarator equilibrium Galeev et al. (1969), this work will consider the simpler problem of an axisymmetric tokamak equilibrium. Our motivation for considering this simpler problem is to investigate the potential for evaluating metrics of energetic particle transport via the solution to an inhomogeneous adjoint problem. As described below, the inhomogeneous adjoint of the drift kinetic equation provides a direct means of computing the average confinement time of an ion or electron as a function of the particle’s initial phase space position Särkimäki (2020), a quantity we will refer to as the mean escape time. While this adjoint problem can be solved by a variety of numerical methods, we will consider a physics-informed neural network (PINN) Raissi et al. (2019) in the present work. Physics-informed neural networks correspond to a machine learning approach capable of integrating both physics and data into the training of a neural network. A distinguishing feature of a PINN is that it can be extended to learn the parametric dependence of the solution of a PDE Sun et al. (2020). Noting that optimizing the transport properties of energetic particle orbits often requires exploring a broad range of magnetic field configurations, obtaining a rapid surrogate capable of computing quantities of interest such as the mean escape time of energetic particles provides a powerful tool for assessing energetic particle transport.
@@
Our aim in the present paper is not to provide such a parametric solution, but to instead investigate the ability of a PINN to solve the drift kinetic equation for an energetic ion population. The extension to a parametric solution will be treated in a future work. Specifically, while PINNs have previously been used to solve the relativistic Arnaud et al. (2024); McDevitt et al. (2025); Arnaud et al. (2025) and non-relativistic McDevitt and Tang (2024) Fokker-Planck equations, the large time scale separation between the rapid bounce or transit times and slow collision time of energetic particles in tokamak geometry, represents a far more challenging problem. Utilizing advanced optimization algorithms Urbán et al. (2025); Kiyani et al. (2025); Wang et al. (2025), we aim to show that PINNs are able to robustly resolve several features of this challenging system, even in the absence of data. As described further below, as the energy of the ions is increased, leading to a greater time scale separation between the fast bounce/transit motion and slow collision time of energetic ions, our implementation of a PINN struggles to quantitatively describe the escape time of the best confined ion orbits, although it does resolve the solution in the outer region of the tokamak including phase space regions of prompt ion loss Hinton and Chu (1985).
@@
As an alternative means of evaluating the fast ion mean escape time, we introduce a new drift-kinetic solver based on the JAX framework. While in the present work this solver is only used to verify solutions from the PINN, future work will leverage particle data generated by this drift-kinetic solver to help accurately resolve the mean escape time of well confined ion orbits, along with facilitating the training of the PINN over a broad range of plasma parameters. This extended PINN framework, trained using both physics and particle data, would thus provide an efficient surrogate for fast ion confinement appropriate for use inside optimization loops.
@@
The remainder of this paper is organized as follows. Section II derives an adjoint problem for the mean escape time of an energetic particle. Brief overviews of the physics-informed deep learning framework and JAX particle-based solver are given in Sec. III. The mean escape time of an energetic ion species is computed in Sec. IV. Conclusions and a discussion of future directions are given in Sec. V
@@ II Adjoint of Steady State Drift Kinetic Equation
In this section we will define an adjoint problem based on the inhomogeneous steady state drift kinetic equation. Our starting point will be to define the Green’s function $F_{s}\left(\mathbf{z};\mathbf{z}_{0}\right)$ for the steady state drift kinetic equation Helander and Sigmar (2002):
@@
$C_{s}\left(F_{s}\right)=\frac{\nu_{D}}{2}\frac{\partial}{\partial\xi}\left[\left(1-\xi^{2}\right)\frac{\partial F}{\partial\xi}\right],$ (2)
@@
where $r_{wall}$ is the location of the surrounding wall, taken as the plasma minor radius $a$, for simplicity. With these boundary conditions particles are not allowed to enter the spatial domain, but particles are able to exit. The resulting Green’s function can thus be physically interpreted as the steady state particle distribution resulting from a unit source located at $\mathbf{z}=\mathbf{z}_{0}$.
To derive an adjoint problem for the mean escape time of an energetic particle, we begin by multiplying the left hand side of Eq. (1a) by the solution to the adjoint equation $T_{s}$, which after successive integrations by parts, yields:
@@
$=\int d^{3}vd^{3}xF_{s}\left[-\mathbf{\dot{X}}\cdot\nabla T_{s}-\dot{V}_{\|}\frac{\partial T_{s}}{\partial v_{\|}}-C^{*}_{s}\left(T_{s}\right)\right]$
$+\int d^{3}v\int d\Gamma\left.\left[U_{r}T_{s}F_{s}\right]\right|_{r=r_{wall}}.$ (5)
@@
An adjoint problem for the mean escape time Grasman et al. (2013); Liu et al. (2017) of an energetic particle can be derived by considering the inhomogeneous adjoint equation
@@
Substituting Eq. (6) along with the steady state Green’s function defined by Eq. (1a) into the right and left hand sides, respectively, of Eq. (5), yields
@@
Further noting that $N_{s}$ will scale directly with the confinement time of the particles, $T_{s}\left(\mathbf{x}_{0},\mathbf{v}_{0}\right)$ will act as a measure of how well particles injected at a given phase space location are confined on average, a quantity often referred to as the mean escape time or expected exit time Grasman et al. (2013). We note that for the present example, all ions will escape for $t\to\infty$, such that the mean escape time will be finite throughout the domain.
@@ III Physics-Constrained Deep Learning Framework
The primary focus of the remainder of this work will be the solution of the adjoint problem derived in Sec. II. A brief overview of our implementation of a PINN is given in Sec. III.1, where the reader is referred to Refs. Cuomo et al. (2022); Wang et al. (2023); Toscano et al. (2025) for additional information on this rapidly expanding area. In addition, a newly developed GPU accelerated particle-based solver for the drift kinetic equation is described in Sec. III.2. For the present work, this particle-based solver will be used to validate the PINN’s solutions to the drift kinetic equation.
@@ III.1 Physics-Informed Neural Networks
Physics-constrained machine learning methods Karniadakis et al. (2021) seek to incorporate physical constraints into the training of a neural network. Physics-informed neural networks correspond to a prominent example of this area. An attractive feature of PINNs is that they can be used to solve PDEs and ODEs directly in the absence of data van Milligen et al. (1995); Raissi et al. (2019). They also provide a mechanism through which physical constraints can be used to enrich sparse data sets Cai et al. (2021); Mathews et al. (2022); McDevitt et al. (2024). In its simplest form, the loss function of a PINN can be written as:
@@
where $T$ is the quantity we seek to predict, the mean escape time for the current work, $\mathcal{R}$ is the residual of the governing PDE, $w_{PDE}$ is a scalar weight applied to the PDE bias of the loss, $\mathbf{z}=\left(\mathbf{x},\mathbf{v}\right)$ are the independent variables, and $\bm{\lambda}$ describes physical parameters, such as collisionality or inverse aspect ratio. The first term represents a bias due to data, where the data may be boundary conditions, or synthetic or experimental data. The second term is a bias due to the governing equations, which for the present work will be the inhomogeneous adjoint of the steady state drift kinetic equation. In the absence of synthetic or experimental data, the first term will only contain the boundary conditions. Successfully minimizing the loss will thus correspond to identifying a solution to the PDE while satisfying the boundary conditions. Unlike traditional PDE solvers, a parametric solution to a PDE can be straightforwardly obtained by training across a broad range of physics parameters $\bm{\lambda}$. Thus, while the training time of such a model may be substantially longer than a traditional PDE solver, once trained, the fast inference time of a PINN allows for fast online prediction times across a broad range of plasma conditions. This property can be used for the development of fast surrogate models, where some recent examples include runaway electron generation McDevitt (2023); Arnaud et al. (2025), MHD equilibrium Jang et al. (2024), plasma thrusters Luo et al. (2025), or non-thermal ion distributions McDevitt and Tang (2024).
Key elements in obtaining accurate results from PINNs correspond to (i) the identification of an appropriate neural network architecture, (ii) the phase space weighting of the residual, and (iii) the selection of an efficient optimization algorithm for minimizing the loss. Regarding item (i) we will employ a fully connected feedforward neural network, but with an ‘output layer’ that enforces positivity as a hard constraint. In so doing, this both prevents certain unphysical solutions from being predicted by the PINN, and perhaps more importantly, narrows the space of solutions that the optimizer searches across. Defining $T_{NN}$ to be the output of the hidden layers of the neural network, a simple transformation that enforces positivity is given by
@@
In addition to enforcing positivity, the exponential dependence of $T_{s}$ on the output of the hidden layers $T_{NN}$ provides a natural means of capturing the large disparity in escape times for energetic ions. Specifically, ions born near the plasma core are expected to be well confined, where they will only be lost by the relatively slow process of collisional transport, in contrast to those born near the edge, which may be lost on a very short time scale by direct orbit loss.
@@
$\text{Loss}=\frac{1}{N_{bdy}}\sum^{N_{bdy}}_{i=1}\left[T_{s,i}-T_{s}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\right]^{2}+\frac{w_{pde}}{N_{PDE}}\sum^{N_{PDE}}_{i=1}\left[\mathcal{F}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\mathcal{R}\left(\mathbf{z}_{i};\bm{\lambda}_{i}\right)\right]^{2}.$ (13)
@@
Here, the boundary conditions defined by Eq. (8) are enforced by the first term, $\mathcal{R}$ is the residual of the inhomogeneous adjoint defined by Eq. (6), and the choice of $A$ sets how much emphasis is given to the edge region. In particular, noting that near the edge $T_{s}$ will be of the order of a transit time, since ions on loss orbits will directly escape from the plasma, if $A$ is chosen to satisfy $A\ll 1$, Eq. (12) will scale as $\propto 1/T_{s}$, such that the edge region will be heavily weighted, but the inner region of the plasma with $T_{s}\gg 1$, will receive minimal weighting. We thus expect this choice to allow for the outer region of the plasma to be accurately described, at the expense of poor accuracy in the inner region. In contrast, by choosing a larger value of $A$, improved accuracy for the inner region can be achieved, at the expense of having a looser connection to the edge boundary condition, which is vital for achieving a physically meaningful solution. We have found that values of $A$ between 10 and 100 allow for robust convergence of the PINN, where all results in the present paper are for $A=100$. In addition, the relative weight of the PDE residual compared to the boundary condition is set by the scalar $w_{PDE}$ weighing the first term in Eq. (13). We have found $w_{PDE}=100$ provides relatively rapid convergence of the loss.
@@
The last key component of our PINN implementation will be the choice of an effective optimization algorithm. While the ADAM optimizer Kingma and Ba (2014) is used ubiquitously across a range of machine learning applications, this first order optimization algorithm often is not able to achieve low losses for challenging PDEs. A common strategy is thus to use ADAM for the first phase of training, after which limited memory BFGS (L-BFGS) Liu and Nocedal (1989), a second order optimization algorithm, is used to more tightly converge the loss. Motivated by Refs. Wang et al. (2025) and Urbán et al. (2025); Kiyani et al. (2025), we have adopted a different optimization strategy that we have found to yield substantially improved accuracy. The first optimizer that we apply corresponds to the quasi-second order optimizer SOAP Vyas et al. (2024). This scalable optimizer exhibits similar computational performance as ADAM, but can yield losses up to an order of magnitude smaller than those achievable by ADAM across a broad range of challenging applications of PINNs Wang et al. (2025). For the second stage of optimization, we utilize the Self-Scaled Broyden (SSBroyden) method Al-Baali et al. (2014). Like L-BFGS, SSBroyden approximates the Hessian, but introduces additional degrees of freedom into the optimization algorithm that have yielded substantial improvements in training across a range of complex PDEs Urbán et al. (2025); Kiyani et al. (2025). While each training epoch is substantially more computationally demanding compared to ADAM or SOAP for the SciPy implementation we are currently employing, we have found it results in substantially faster convergence of the loss for the adjoint problem treated in this work. The primary limitations of SSBroyden as an optimization routine is that the SciPy implementation used in the present work does not efficiently utilize GPUs or support mini-batching. These limitations substantially increase the training time of the SSBroyden phase of training, and limit the number of training points employed.